OpenStudy (anonymous):

Cindy borrowed the same amount of money from a bank and a friend. The table below shows the amount, in dollars, that Cindy would owe them after different numbers of years: Year 1 2 3 4 Bank 102.50 105.06 107.69 110.38 Friend 103 106 109 112 Which statement is true about the money Cindy would owe the bank and her friend after 30 years? She would owe the bank more money. She would owe her friend more money. She would owe both the same amount of money. She would owe her friend twice the amount she borrowed.

OpenStudy (anonymous):

rate of interest of bank should be known

OpenStudy (anonymous):

@Amber1997

OpenStudy (anonymous):

so D? @swarup169

OpenStudy (anonymous):

@Amber1997

OpenStudy (anonymous):

the interest is increasing by 0.06 each year :P

OpenStudy (anonymous):

so D? @aajugdar

OpenStudy (anonymous):

calculate it xD

OpenStudy (anonymous):

ok i will tell you how

OpenStudy (anonymous):

first year's interest is 2.56

OpenStudy (anonymous):

B

OpenStudy (anonymous):

??

OpenStudy (anonymous):

2nd year's 2.63 and third year's 2.69 now the difference btwn them is 0.06 now to find what it will be after thirty years 2.63+(30-1)0.06 = 4.3 now the sum will be

OpenStudy (anonymous):

30(2.56+4.3)/2 =102

OpenStudy (anonymous):

@kirbykirby

OpenStudy (anonymous):

so final amount she has to give to bank is 102.50+ 102 =204.50

OpenStudy (anonymous):

now as for her friend it will be 30*3 =90 103+90= 193

OpenStudy (anonymous):

so she owes bank more money :P

OpenStudy (anonymous):

so D?

OpenStudy (anonymous):

its A :O

OpenStudy (anonymous):

what did i solve just nw -.-

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

last question!

OpenStudy (anonymous):

m checking

OpenStudy (anonymous):

hey A is correct :)

OpenStudy (anonymous):

for this one too

OpenStudy (anonymous):

did you get how

OpenStudy (anonymous):

OS is laggy sigh